# A Survey of Reinforcement Learning For Economics

### 23 Mar 2026 v4

@@ 6 $\sim 10^{8}$ Failed
Where classical solutions exist, RL underperforms: the base-stock policy remains highly competitive when properly calibrated, and DRL required millions of training transitions to approach performance levels achievable through closed-form calculation. RL struggles particularly with multi-echelon coupling, where upstream orders affect downstream costs many periods later.142142142Credit assignment refers to the difficulty of determining which past actions caused a delayed reward or cost. In multi-echelon systems, an upstream ordering decision today may not affect customer-facing costs for many periods, making it difficult for RL to learn the causal connection. The value proposition for RL lies in problems where analytical solutions are unavailable: non-stationary demand, complex operational constraints, or cost structures that do not admit tractable decomposition.143143143Residual RL (Silver et al., 2018b) offers a middle ground: rather than replacing classical policies, learn a correction $\pi_{\mathrm{final}}(s)=\pi_{\mathrm{classical}}(s)+\pi_{\mathrm{learned}}(s)$, preserving the classical solution’s performance while allowing RL to handle constraints or non-stationarity the analytical method cannot. This connects to perturbation methods in economics, where one linearizes around a known solution and solves for deviations. See Ibarz et al. (2021) for a survey of residual and demonstration-augmented RL in robotics. Even in these settings, successful deployment remains rare and requires extensive simulation infrastructure, domain expertise, and careful calibration against classical baselines.
@@ 7.6 Real-Time Bidding
where $\mathrm{cvr}_{t}\in\{0,1\}$ is the conversion indicator. Budget evolves as $B_{t+1}=B_{t}-\mathbbm{1}\{b_{t}\geq c_{t}\}\cdot c_{t}$, and the episode terminates when $B_{t}\leq 0$ or t=0t=0.144144144The state space is discretized into budget bins and time bins. Wu et al. (2018) augment the state with bid landscape features derived from historical clearing price distributions, giving the agent information about current market competitiveness.
@@ 7.7 Simulation Study: Bus Engine Replacement
I conclude with a simulation demonstrating that RL matches dynamic programming on a classical economics benchmark. The bus engine replacement problem, introduced by Rust (1987), models a fleet manager’s monthly decision whether to replace engines based on accumulated mileage. Replacement incurs a fixed cost but resets mileage to zero; continued operation incurs maintenance costs increasing in mileage.
I extend the single-engine problem to a fleet of $N$ engines with a capacity constraint limiting replacements per period. The state $s=(m_{1},\ldots,m_{N})$ records discretized mileage for each engine. Actions are subsets of engines to replace, subject to the capacity constraint. The per-period cost is $c(s,a)=\alpha\sum_{i}m_{i}+\beta|a|$, where $\alpha$ is the operating cost per unit mileage and $\beta$ is the replacement cost.145145145The mileage-dependent operating cost $\alpha\sum_{i}m_{i}$ follows Rust’s original specification $c(x,\theta_{1})=\theta_{11}x$, creating a non-trivial threshold replacement policy. The fleet extension uses deterministic mileage increments to isolate the combinatorial scaling challenge that arises from the joint state of multiple engines. Mileage evolves deterministically; replaced engines reset to $m=0$; others increment by one bin.146146146With $M=6$ mileage bins, the state space is $6^{N}$: 1,296 states at $N=4$, 7,776 at $N=5$, 46,656 at $N=6$. Value iteration is feasible for $N\leq 5$.
@@ 8 Structural Estimation with Reinforcement Learning
Several recent papers have used RL training loops,147147147Throughout this chapter, the RL training loop runs entirely inside the econometrician’s computational model; the agent never interacts with real economic agents or markets but serves as a numerical method for solving the Bellman equation within a structural estimation procedure (Section 2). There is no execution phase in the usual sense. namely Q-learning, temporal-difference learning, policy gradient, and actor-critic methods, to solve structural economic models at scales where conventional dynamic programming fails.148148148I exclude papers that use neural network function approximation without an RL training mechanism. Inverse reinforcement learning is treated in the sister survey (Rust and Rawat, 2026). Throughout this chapter I adopt a unified notation. An MDP is a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition kernel, $r(s,a)$ is the per-period reward, and $\gamma\in[0,1)$ is the discount factor.149149149Several of the papers reviewed here use $\beta$ for the discount factor, following economics convention. I translate all results to $\gamma$ for consistency with the RL literature and the rest of this survey. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions. The value function under $\pi$ is $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s\right]$, and the action-value function is $Q^{\pi}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}(s^{\prime})]$. The optimal value function satisfies $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)$.
@@ 8.1.1 TD Learning for CCP Estimation
The first is the linear semi-gradient method.150150150The method is called “semi-gradient” because it computes the gradient of only the prediction $\phi(a,s)^{\top}w$ with respect to $w$, not the full TD error including the bootstrap target $\phi(a^{\prime},s^{\prime})^{\top}w$. This avoids differentiating through the target but sacrifices guaranteed convergence in some settings. Approximate $h(a,s)\approx\phi(a,s)^{\top}w$ where $\phi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{p}$ is a vector of basis functions (e.g., polynomials in state variables) and $w\in\mathbb{R}^{p}$ are the weights to be estimated. The TD(0) fixed-point equation is
@@
and then fits $\hat{h}^{(k)}$ by regressing $Y_{it}^{(k)}$ on $(a_{it},s_{it})$ using any machine learning method, including LASSO, random forests, or neural networks.151151151LASSO (Tibshirani, 1996) adds an $\ell_{1}$ penalty to a regression loss, shrinking many coefficients to zero for sparse solutions; random forests average many decision trees fit to random subsamples and feature subsets for nonparametric estimation; neural networks compose affine transformations with elementwise nonlinearities across multiple layers. This is the first DDC estimator compatible with arbitrary ML prediction methods, enabling application to very high-dimensional state spaces.
With $\hat{h}$ and $\hat{g}$ in hand, structural parameters are recovered by pseudo-maximum likelihood estimation (PMLE). For continuous state spaces, Adusumilli et al. (2022) derive a locally robust correction to the PMLE criterion that accounts for the nonparametric first-stage estimation of value terms, restoring $\sqrt{n}$-convergence of $\hat{\theta}$.152152152An estimator has $\sqrt{n}$-convergence if its error shrinks at rate $n^{-1/2}$ where $n$ is the sample size. This is the standard parametric rate; slower rates (e.g., $n^{-1/4}$) indicate efficiency loss from nonparametric first stages. The PMLE score is $m(a,s;\theta,h,g)=\partial_{\theta}\ln\pi(a,s;\theta,h,g)$, where $\pi$ is the logit choice probability with continuation value $V(a,s)=h(a,s)^{\top}\theta+g(a,s)$. The naive estimator solves $\mathbb{E}_{n}[m(a,s;\theta,\hat{h},\hat{g})]=0$, but with continuous states this moment condition is not orthogonal to the first-stage estimates and converges slower than $\sqrt{n}$. The locally robust moment adds a debiasing correction:
@@
where $\lambda(a,s;\theta)$ solves a backward recursion that propagates the influence of estimation error through the dynamic structure.153153153The term in braces is the temporal-difference error of the continuation value $V$. The adjoint $\lambda$ weights this TD error by its marginal impact on the PMLE score. See Online Appendix B.3 of Adusumilli et al. (2022) for the derivation. The corrected estimator $\hat{\theta}_{LR}$ solves $\mathbb{E}_{n}[\zeta_{n}]=0$ and is computationally no harder than the naive PMLE, since the correction is constant in $\theta$.
@@
Under regularity conditions, the linear semi-gradient estimator $\hat{h}$ satisfies $\|\hat{h}-h\|_{2}=O_{P}(n^{-1/2}(T-1)^{-1/2})$, where $\|\cdot\|_{2}$ denotes the $L^{2}(P)$ norm.154154154The $L^{2}(P)$ norm is $\|f\|_{2}=(\int f(x)^{2}\,dP(x))^{1/2}$, measuring average squared deviation under the probability measure $P$. This is the natural norm for mean-squared-error analysis.
@@ 8.1.2 Policy Gradient for DDC Estimation
The outer loop is SMM.155155155SMM (Gourieroux et al., 1993) estimates structural parameters by matching moments from simulated data to moments from observed data, avoiding direct evaluation of the likelihood function. Define a vector of data moments $\mathbf{M}_{d}$ computed from the observed panel. For candidate structural parameters $\theta$ and transition parameters $\xi$, simulate the model to produce simulated moments $\mathbf{M}_{s}(\theta,\xi)$. The estimator minimizes
@@
The inner loop parametrizes the choice probability directly as a logistic function of state variables. For a binary choice $J_{t}\in\{0,1\}$, the general form is $\Pr(J_{t}=1\mid\boldsymbol{X}_{t};\boldsymbol{\gamma}(\theta))=\text{logistic}(\boldsymbol{X}_{t}\boldsymbol{\gamma}(\theta))$, where $\boldsymbol{\gamma}(\theta)$ are policy parameters that depend on the structural parameters.156156156The linear index can be replaced by higher-order terms of $\boldsymbol{X}_{t}$ or deep neural networks; the method requires only that the gradient $\nabla_{\boldsymbol{\gamma}}\log\pi_{\boldsymbol{\gamma}}$ has a closed form. In their application to a Rust bus engine model with unobserved bus condition $S_{t}^{*}$, this takes the form
@@ 8.2.2 TD Learning for Merger Analysis with Innovation
The value function update for firm $i$ is159159159This stochastic approximation update, introduced by Pakes and McGuire (1994) for dynamic oligopoly computation, is closely related to temporal-difference learning in the RL literature. The original algorithm uses visit-count averaging ($\alpha_{k}=1/k$ where $k$ counts visits to each state) rather than a fixed learning rate.
@@
where $\Pi_{i}(\Omega,\mathbf{a})$ denotes firm $i$’s per-period profit given industry state $\Omega$ and the joint action profile $\mathbf{a}$ (investment, entry/exit, merger decisions),160160160I use $\Pi_{i}$ for firm $i$’s per-period profit to avoid confusion with policy $\pi$. $\gamma$ is the discount factor, and $\Omega^{\prime}$ is the realized next-period state. The algorithm uses $\varepsilon$-decreasing exploration to prevent convergence to locally suboptimal equilibria.
The equilibrium computation follows the Pakes-McGuire iterative scheme.161161161The Pakes-McGuire algorithm (Pakes and McGuire, 1994) computes Markov perfect equilibria by iterating: (1) given current value functions, compute best-response strategies; (2) given strategies, update value functions via simulation. Convergence is not guaranteed but works well in practice. The algorithm simulates long industry histories, updates each firm’s value function via (83), re-derives best-response strategies from updated values, and repeats.
@@
Two related papers merit brief mention. Lomys and Magnolfi (2024) develop structural estimation methods for strategic settings where agents use learning algorithms (specifically regret-minimizing rules) rather than playing a fixed equilibrium. They impose an “asymptotic no-regret” condition as a minimal rationality requirement and derive identification results for payoff parameters. Covarrubias (2022) uses deep RL to study oligopolistic pricing in a New Keynesian framework, representing firms’ pricing policies as neural networks $\pi_{\phi}(a|s)$. The method uncovers multiple equilibria ranging from competitive to collusive pricing.162162162The most prominent example of algorithmic collusion is Calvano et al. (2020), who showed that independent Q-learning agents in a repeated Bertrand pricing game learn to sustain supra-competitive prices and punish deviators without explicit communication. This result and its implications for competition policy are treated in the companion thesis chapter (Rawat, 2026).
@@ 8.3.1 RL for Sequential Price Mechanisms
Brero et al. (2021) use RL to design optimal sequential price mechanisms (SPMs), a class of indirect auction mechanisms where agents are approached in sequence and offered menus of items at posted prices. SPMs generalize both serial dictatorship and posted-price mechanisms and essentially characterize all strongly obviously strategyproof (SOSP) mechanisms (Pycia and Troyan, 2023).163163163A mechanism is strategyproof if truthful reporting is a dominant strategy. It is obviously strategyproof if this dominance is apparent even to boundedly rational agents; strongly obviously strategyproof (SOSP) adds that dominance holds at every information set (Pycia and Troyan, 2023).
The mechanism design problem is formulated as a partially observable Markov decision process (POMDP).164164164In a POMDP, the agent cannot directly observe the full state. It maintains a belief distribution over possible states, updated via Bayes’ rule as new observations arrive. This captures the mechanism designer’s uncertainty about bidder valuations. The state at round $t$ includes the set of remaining items $\rho_{t}^{\text{items}}\subseteq[m]$, remaining agents $\rho_{t}^{\text{agents}}\subseteq[n]$, the partial allocation $\mathbf{x}_{t}$, and the agents’ (unobserved) valuation functions. The action $a_{t}=(i_{t},\{p_{j}^{t}\}_{j\in\rho_{t-1}^{\text{items}}})$ specifies which agent to visit next and what prices to offer. The observation is the agent’s purchase decision, from which the mechanism can update beliefs about valuations. The reward is the objective function evaluated at the final allocation:
@@
The policy maps from a sufficient statistic of the observation history to actions. Brero et al. (2021) show that this statistic can be represented compactly. The set of remaining items and agents suffices for independent valuations, while the full allocation matrix is needed for correlated valuations. They train the mechanism policy using Proximal Policy Optimization (PPO),165165165Proximal Policy Optimization (Schulman et al., 2017) is a policy gradient algorithm that constrains each update to a trust region, preventing large destabilizing policy changes. It is described in Chapter 2. which handles the discrete action space (agent selection) and continuous action space (price setting) of the POMDP.

### 24 Mar 2026 v5

@@ 6 $\sim 10^{8}$ Failed
Where classical solutions exist, RL underperforms: the base-stock policy remains highly competitive when properly calibrated, and DRL required millions of training transitions to approach performance levels achievable through closed-form calculation. RL struggles particularly with multi-echelon coupling, where upstream orders affect downstream costs many periods later.144144144Credit assignment refers to the difficulty of determining which past actions caused a delayed reward or cost. In multi-echelon systems, an upstream ordering decision today may not affect customer-facing costs for many periods, making it difficult for RL to learn the causal connection. The value proposition for RL lies in problems where analytical solutions are unavailable: non-stationary demand, complex operational constraints, or cost structures that do not admit tractable decomposition.145145145Residual RL (Silver et al., 2018b) offers a middle ground: rather than replacing classical policies, learn a correction $\pi_{\mathrm{final}}(s)=\pi_{\mathrm{classical}}(s)+\pi_{\mathrm{learned}}(s)$, preserving the classical solution’s performance while allowing RL to handle constraints or non-stationarity the analytical method cannot. This connects to perturbation methods in applied mathematics, where one linearizes around a known solution and solves for deviations. See Ibarz et al. (2021) for a survey of residual and demonstration-augmented RL in robotics. Even in these settings, successful deployment remains rare and requires extensive simulation infrastructure, domain expertise, and careful calibration against classical baselines.
@@ 7.6 Real-Time Bidding
where $\mathrm{cvr}_{t}\in\{0,1\}$ is the conversion indicator. Budget evolves as $B_{t+1}=B_{t}-\mathbbm{1}\{b_{t}\geq c_{t}\}\cdot c_{t}$, and the episode terminates when $B_{t}\leq 0$ or t=0t=0.146146146The state space is discretized into budget bins and time bins. Wu et al. (2018) augment the state with bid landscape features derived from historical clearing price distributions, giving the agent information about current market competitiveness.
@@ 7.7 Simulation Study: Bus Engine Replacement
I conclude with a simulation demonstrating that RL matches dynamic programming on a classical benchmark. The bus engine replacement problem, introduced by Rust (1987), models a fleet manager’s monthly decision whether to replace engines based on accumulated mileage. Replacement incurs a fixed cost but resets mileage to zero; continued operation incurs maintenance costs increasing in mileage.
I extend the single-engine problem to a fleet of $N$ engines with a capacity constraint limiting replacements per period. The state $s=(m_{1},\ldots,m_{N})$ records discretized mileage for each engine. Actions are subsets of engines to replace, subject to the capacity constraint. The per-period cost is $c(s,a)=\alpha\sum_{i}m_{i}+\beta|a|$, where $\alpha$ is the operating cost per unit mileage and $\beta$ is the replacement cost.147147147The mileage-dependent operating cost $\alpha\sum_{i}m_{i}$ follows Rust’s original specification $c(x,\theta_{1})=\theta_{11}x$, creating a non-trivial threshold replacement policy. The fleet extension uses deterministic mileage increments to isolate the combinatorial scaling challenge that arises from the joint state of multiple engines. Mileage evolves deterministically; replaced engines reset to $m=0$; others increment by one bin.148148148With $M=6$ mileage bins, the state space is $6^{N}$: 1,296 states at $N=4$, 7,776 at $N=5$, 46,656 at $N=6$. Value iteration is feasible for $N\leq 5$.
@@ 8 Structural Estimation with Reinforcement Learning
Several recent papers have used RL training loops,149149149Throughout this chapter, the RL training loop runs entirely inside the econometrician’s computational model; the agent never interacts with real economic agents or markets but serves as a numerical method for solving the Bellman equation within a structural estimation procedure (Section 2). There is no execution phase in the usual sense. namely Q-learning, temporal-difference learning, policy gradient, and actor-critic methods, to solve structural economic models at scales where conventional dynamic programming fails.150150150I exclude papers that use neural network function approximation without an RL training mechanism. Inverse reinforcement learning is treated in the sister survey (Rust and Rawat, 2026). Throughout this chapter I adopt a unified notation. An MDP is a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition kernel, $r(s,a)$ is the per-period reward, and $\gamma\in[0,1)$ is the discount factor.151151151Several of the papers reviewed here use $\beta$ for the discount factor, following economics convention. I translate all results to $\gamma$ for consistency with the RL literature and the rest of this survey. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions. The value function under $\pi$ is $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s\right]$, and the action-value function is $Q^{\pi}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}(s^{\prime})]$. The optimal value function satisfies $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)$.
@@ 8.1.1 TD Learning for CCP Estimation
The first is the linear semi-gradient method.152152152The method is called “semi-gradient” because it computes the gradient of only the prediction $\phi(a,s)^{\top}w$ with respect to $w$, not the full TD error including the bootstrap target $\phi(a^{\prime},s^{\prime})^{\top}w$. This avoids differentiating through the target but sacrifices guaranteed convergence in some settings. Approximate $h(a,s)\approx\phi(a,s)^{\top}w$ where $\phi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{p}$ is a vector of basis functions (e.g., polynomials in state variables) and $w\in\mathbb{R}^{p}$ are the weights to be estimated. The TD(0) fixed-point equation is
@@
and then fits $\hat{h}^{(k)}$ by regressing $Y_{it}^{(k)}$ on $(a_{it},s_{it})$ using any machine learning method, including LASSO, random forests, or neural networks.153153153LASSO (Tibshirani, 1996) adds an $\ell_{1}$ penalty to a regression loss, shrinking many coefficients to zero for sparse solutions; random forests average many decision trees fit to random subsamples and feature subsets for nonparametric estimation; neural networks compose affine transformations with elementwise nonlinearities across multiple layers. This is the first DDC estimator compatible with arbitrary ML prediction methods, enabling application to very high-dimensional state spaces.
With $\hat{h}$ and $\hat{g}$ in hand, structural parameters are recovered by pseudo-maximum likelihood estimation (PMLE). For continuous state spaces, Adusumilli et al. (2022) derive a locally robust correction to the PMLE criterion that accounts for the nonparametric first-stage estimation of value terms, restoring $\sqrt{n}$-convergence of $\hat{\theta}$.154154154An estimator has $\sqrt{n}$-convergence if its error shrinks at rate $n^{2}$ where $n$ is the sample size. This is the standard parametric rate; slower rates (e.g., $n^{4}$) indicate efficiency loss from nonparametric first stages. The PMLE score is $m(a,s;\theta,h,g)=\partial_{\theta}\ln\pi(a,s;\theta,h,g)$, where $\pi$ is the logit choice probability with continuation value $V(a,s)=h(a,s)^{\top}\theta+g(a,s)$. The naive estimator solves $\mathbb{E}_{n}[m(a,s;\theta,\hat{h},\hat{g})]=0$, but with continuous states this moment condition is not orthogonal to the first-stage estimates and converges slower than $\sqrt{n}$. The locally robust moment adds a debiasing correction:
@@
where $\lambda(a,s;\theta)$ solves a backward recursion that propagates the influence of estimation error through the dynamic structure.155155155The term in braces is the temporal-difference error of the continuation value $V$. The adjoint $\lambda$ weights this TD error by its marginal impact on the PMLE score. See Online Appendix B.3 of Adusumilli et al. (2022) for the derivation. The corrected estimator $\hat{\theta}_{LR}$ solves $\mathbb{E}_{n}[\zeta_{n}]=0$ and is computationally no harder than the naive PMLE, since the correction is constant in $\theta$.
@@
Under regularity conditions, the linear semi-gradient estimator $\hat{h}$ satisfies $\|\hat{h}-h\|_{2}=O_{P}(n^{2}(T-1)^{2})$, where $\|\cdot\|_{2}$ denotes the $L^{2}(P)$ norm.156156156The $L^{2}(P)$ norm is $\|f\|_{2}=(\int f(x)^{2}\,dP(x))^{1/2}$, measuring average squared deviation under the probability measure $P$. This is the natural norm for mean-squared-error analysis.
@@ 8.1.2 Policy Gradient for DDC Estimation
The outer loop is SMM.157157157SMM (Gourieroux et al., 1993) estimates structural parameters by matching moments from simulated data to moments from observed data, avoiding direct evaluation of the likelihood function. Define a vector of data moments $\mathbf{M}_{d}$ computed from the observed panel. For candidate structural parameters $\theta$ and transition parameters $\xi$, simulate the model to produce simulated moments $\mathbf{M}_{s}(\theta,\xi)$. The estimator minimizes
@@
The inner loop parametrizes the choice probability directly as a logistic function of state variables. For a binary choice $J_{t}\in\{0,1\}$, the general form is $\Pr(J_{t}=1\mid\boldsymbol{X}_{t};\boldsymbol{\gamma}(\theta))=\text{logistic}(\boldsymbol{X}_{t}\boldsymbol{\gamma}(\theta))$, where $\boldsymbol{\gamma}(\theta)$ are policy parameters that depend on the structural parameters.158158158The linear index can be replaced by higher-order terms of $\boldsymbol{X}_{t}$ or deep neural networks; the method requires only that the gradient $\nabla_{\boldsymbol{\gamma}}\log\pi_{\boldsymbol{\gamma}}$ has a closed form. In their application to a Rust bus engine model with unobserved bus condition $S_{t}^{*}$, this takes the form
@@ 8.2.2 TD Learning for Merger Analysis with Innovation
The value function update for firm $i$ is161161161This stochastic approximation update, introduced by Pakes and McGuire (1994) for dynamic oligopoly computation, is closely related to temporal-difference learning in the RL literature. The original algorithm uses visit-count averaging ($\alpha_{k}=1/k$ where $k$ counts visits to each state) rather than a fixed learning rate.
@@
where $\Pi_{i}(\Omega,\mathbf{a})$ denotes firm $i$’s per-period profit given industry state $\Omega$ and the joint action profile $\mathbf{a}$ (investment, entry/exit, merger decisions),162162162I use $\Pi_{i}$ for firm $i$’s per-period profit to avoid confusion with policy $\pi$. $\gamma$ is the discount factor, and $\Omega^{\prime}$ is the realized next-period state. The algorithm uses $\varepsilon$-decreasing exploration to prevent convergence to locally suboptimal equilibria.
The equilibrium computation follows the Pakes-McGuire iterative scheme.163163163The Pakes-McGuire algorithm (Pakes and McGuire, 1994) computes Markov perfect equilibria by iterating: (1) given current value functions, compute best-response strategies; (2) given strategies, update value functions via simulation. Convergence is not guaranteed but works well in practice. The algorithm simulates long industry histories, updates each firm’s value function via (83), re-derives best-response strategies from updated values, and repeats.
@@
Two related papers merit brief mention. Lomys and Magnolfi (2024) develop structural estimation methods for strategic settings where agents use learning algorithms (specifically regret-minimizing rules) rather than playing a fixed equilibrium. They impose an “asymptotic no-regret” condition as a minimal rationality requirement and derive identification results for payoff parameters. Covarrubias (2022) uses deep RL to study oligopolistic pricing in a New Keynesian framework, representing firms’ pricing policies as neural networks $\pi_{\phi}(a|s)$. The method uncovers multiple equilibria ranging from competitive to collusive pricing.164164164The most prominent example of algorithmic collusion is Calvano et al. (2020), who showed that independent Q-learning agents in a repeated Bertrand pricing game learn to sustain supra-competitive prices and punish deviators without explicit communication. This result and its implications for competition policy are treated in the companion thesis chapter (Rawat, 2026).
@@ 8.3.1 RL for Sequential Price Mechanisms
Brero et al. (2021) use RL to design optimal sequential price mechanisms (SPMs), a class of indirect auction mechanisms where agents are approached in sequence and offered menus of items at posted prices. SPMs generalize both serial dictatorship and posted-price mechanisms and essentially characterize all strongly obviously strategyproof (SOSP) mechanisms (Pycia and Troyan, 2023).165165165A mechanism is strategyproof if truthful reporting is a dominant strategy. It is obviously strategyproof if this dominance is apparent even to boundedly rational agents; strongly obviously strategyproof (SOSP) adds that dominance holds at every information set (Pycia and Troyan, 2023).
The mechanism design problem is formulated as a partially observable Markov decision process (POMDP).166166166In a POMDP, the agent cannot directly observe the full state. It maintains a belief distribution over possible states, updated via Bayes’ rule as new observations arrive. This captures the mechanism designer’s uncertainty about bidder valuations. The state at round $t$ includes the set of remaining items $\rho_{t}^{\text{items}}\subseteq[m]$, remaining agents $\rho_{t}^{\text{agents}}\subseteq[n]$, the partial allocation $\mathbf{x}_{t}$, and the agents’ (unobserved) valuation functions. The action $a_{t}=(i_{t},\{p_{j}^{t}\}_{j\in\rho_{t-1}^{\text{items}}})$ specifies which agent to visit next and what prices to offer. The observation is the agent’s purchase decision, from which the mechanism can update beliefs about valuations. The reward is the objective function evaluated at the final allocation:
@@
The policy maps from a sufficient statistic of the observation history to actions. Brero et al. (2021) show that this statistic can be represented compactly. The set of remaining items and agents suffices for independent valuations, while the full allocation matrix is needed for correlated valuations. They train the mechanism policy using Proximal Policy Optimization (PPO),167167167Proximal Policy Optimization (Schulman et al., 2017) is a policy gradient algorithm that constrains each update to a trust region, preventing large destabilizing policy changes. It is described in Chapter 2. which handles the discrete action space (agent selection) and continuous action space (price setting) of the POMDP.